AO Preset Testing and Range Balancing: UTC 2011 May 21

Observer(s):D. Miller, K. Boutsia, JC. Guerra (LBTO) A. Rakich, J. Hill, O. Kuhn (Tucson)
Telescope Operator:S. Allanson (LBTO)
Software: N. Cushing, T. Leibold, D. Cox, C. Biddick, M. Delapena, T. Sargent Telescope Support: J. Urban
Instrument Support: T. Shih

Summary

Details

16:27 (afternoon) Power ON W-Unit

Email sent by Doug to LBT and Arcetri people
-----------------------------------

The Power "ON" and presetAO tests are complete
and were successful.

The Power "ON" had no problems.

The first presetAO failed because the Overhead
lights were on and the ccd49 was saturated

The second presetAO failed because I had sent
the incorrect focal station name.

The third preset AO was successful, but the ccd49
black levels were not correct.  The lower half was
all 0

Alfio walked us through setting the black levels
and we found reasonable values.  This was the
problem of thermal changes on startup.

Fourth presetAO was successful and the both
ccd39 and 47 looked good.

All tests successful.

We did find that I could not send a power "OFF" when
the system was in AOPrepared state.  Alfio had me
run an Operate command and then OFF.  This was
successful.  Alfio said this was an error in logic in
the state machine that had been fixed, but apparently
not updated on the mountain.  We now know the work
around and will be fine.

Thanks for the help Alfio and suggestions from Javier.

17:20 tests complete

Starting nighttime log. Times are in MST.

19:26 Slew to WT10_401. Run dofpia, /First at 19:30. Much too bright outside.

19:44 Still too bright...

20:07 R 8907 = Z4, no pupils found on B. After a few iterations, quit with S.

20:11 WT10_401. Star not in field of chip2. Correct pointing models are in place. Slew again to WT10_401. We do not see a bright star even in the full LBC field.

Pointing Diagnosis

03:30 Slew to reset telescope encoders again.

03:35 Preset to Arcturus (not in the 6x8 arcmin field), although we see a glow.

03:40 Stop MCSPU to switch back to old configuration files. (Dan C. had recovered the failed disk with fsck.)

03:43 Slew to reset telescope encoders again.

03:47 Slew to WT10_401 and find the star.

03:54 Slew to field LBCCollimationTemperature_test.ob I did not rangebal since the cursor value was not being picked up, possibly a connection problem.

04:12 Blue has converged, Red is still converging... lots of -ve Z4 corrections sent. Temperatures:
SX: 1.409/1.282/0.5
DX: 1.004/1.283/0.5

04:15 Stopping dofpia.

TCS Test

04:17 Slew to PG1633_gr_4pos

04:24 Exp x30 on same field

04:34 Change to Right authorization only. Disconnect/Reconnect LBC.

04:36 Exp x30 for PG1633 with telescope function only on the Right.

04:44 Change to Left authorization only. Disconnect/Reconnect LBC.

04:45 Exp x30 for PG1633 with telescope function only on the Left.

04:53 Back to BOTH LBCs authorized

04:54 Exp x30 for PG1633

05:04 Back to the LBCCollimationTemperature_test field at 15:30 +70

05:07 DOFPIA
-263 293 -156 105 -68 7 -13 y 1.08 0.17 /newdata/lbcb.20110521.050759.fits V-BESSEL
-945 -134 -5 -19 67 -68 12 y 0.77 0.39 /newdata/lbcr.20110521.050839.fits r-SLOAN

05:12 Turn off LBC. Reconfigure telescope for Bent Gregorian

05:55 ACQUIRE preset to WT10_401

TCS Tests at Bent Gregorian from LBTtools

WT10_401 and BS9156

Can JMH break the 75 sec per command threshold?

SX IE=-100 CA=+74
DX IE=-6 CA=-5

07:00 Tests from the LUCI interface

AO Testing

07:15 - Starting AO test - WFS unit power ON

07:27 - Sending preset AO613 -> iif-preset, AO613, MODE=ACE, RS=0,GS=1

07:31 - preset failed - Arbitrator interface generic error
  • the preset had a command time out
  • form the GCS gui we see that it is collimating

07:47 - preset AO - waiting for it to time out...
  • contacting Alfio - he sais that.the interface between AOS and supervisor changed between BP10 and BP11.
  • either update supervisor (Arbitrator) or switch back to BP10.

07:55 - switching to BP10 mcspu

08:07 - Sending AO613 preset again -> iif-preset, AO613, MODE=ACE, RS=0,GS=1
  • preset failed - GCS could not find the star - pointing IE and CA need to be set

08:09 - Send preset again -> iif-preset, AO613, MODE=ACE, RS=0,GS=1
  • preset failed again - we go to a nearby bright star to correct pointing
  • preset AO was successful though
  • problems with the pointing

08:42 - Still problems with the TCS configuration - shut down the build
  • log out and log in again!

08:52 - Acquire preset - still problems * tertiary was out of position ... now it is ok

09:15 - back on line
  • we see stars in the GCS, but it is out of focus

09:18 - Sending adaptive preset -> iif_preset, AO677, MODE= ACE, RS=0, GS=8,PA=180
  • AO preset - successful
  • Active Optics collimated the telescope in 5 or 6 iteration. A lot of focus and astigmatism

09:24 - Sending a runAO ->iif_runAO (acquireRefAO + startAO)
  • acquireRefAO failed - unspecified idl error
  • the slow closed test loop failed to close. Slopes are still being sent to AdSec and being skipped.

09:28 - again - iif_runAO
  • it is sending slopes and it does not time out
  • trying to stop the slope computer - it is not responding

09:32 - send stopAO
  • the pyramid is still sending slopes, or at least the AdSec says it is receiving slopes and skipping them.

09:36 - turn offf Slope BCU - is not responding - still getting frames. It was responding. This was a GUI update problem. We killed the system process GUI and restarted and the new state (On then Off) of the Slope BCU was shown.

09:37 - Power off W_Unit from the AOS GUI - AdSec is still getting frames and skipping them, even though all W-Unit systems are powered off
  • rested the secondary - we don't get frames anymore

09:40 - flattened the secondary -> flatAO
  • power on W-Unit from the AOS GUI. Success

09:45 - sending adaptive preset -> iif_preset, AO677, MODE= ACE, RS=0, GS=8,PA=180
  • AO preset successful
  • iif_runAO
  • close loop - done!
  • Star seen on the IRTC
  • we did not take images because the IRTC was in large field FOV. We switched to the Small FOV, but the star shifted out of the field of view
  • Adsec started skipping all frames
  • this happened because we did not offload tip/tilt - it started skipping frames - guide star shifted out of the field

09:55 - stop AO - handing over to Andrew

Range Balance Testing

RFBG performance Data

11:28 preset, BS9177, mode=ACTIVE, GS=8 Failed, couldn't find guide star

11:30 Steve preset to pointing star. None Found

11:34 Steve force tertiary to different focal station and then back to RFBG

11:38 Cleared Active Optics and C00. Star visible. Set IE and CA

11:39 preset, BS9177, mode=ACTIVE, GS=8 too dim with sky getting bright

11:39 preset, BS9177, mode=ACTIVE, GS=5, TS=3 Brighter star for Active optics and for IRTC

11:47 dice5, dist=10, InformationTechnologyER=5, EXP=1.0, NUM=5 The sky was very bright, SH images looking like tadpoles, GCS finally stopped send active optic corrections. Done!

11:54 Steve starts closing the Chamber

-- DougMiller - 20 May 2011
Topic revision: r9 - 25 Jul 2014, DouglasFisher
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback